= 211.0 3) A diffuse surface at T = (600 + 200C)K.has the spectral, hemispherical emissivity illustrated. Determine the following
Answer #1

diffuse surface at Temperature, T= (600 + 2004)* G = 21 A >> Tempesatwie, T = 600+200x 21 = 4800K Spect sal hemispherical emi5 2 towie E = (0-4) Ex, bodd 0.8 da + + Elboda D Eb Eь Eь Temperature, T=4800K » E= 0.4 F822,4m) + 0.3 [ Floosum) - * (0-24m)I hope you understand. Comment if you have any doubt.